# VideoNSA: Native Sparse Attention Scales Video Understanding

### v1

@@ 9 The Use of Large Language Models
  • Visualization of attention pattern in each branch in Section G.
  • More results on branch combination in Section H.
  • More results on information scaling study in Section I.
  • More results on attention scaling study in Section J.
  • Full gate values distribution in Section K.
  • More inter-head gate similarites visualization in Section L.
  • More analysis about attention sinks on various sparse attention settings can be found in Section M.
  • Visualization of attention sinks in dense attention in Section N.
@@ A.1 Efficient Video Understanding
Video understanding systems typically convert videos into long sequences of vision tokens, which can easily exceed GPU memory and slow down inference as the video length grows. To address this, existing work mainly address this by token compression, alternative sequence modeling, and KV-cache compression. One important line of work emphasizes token compression. Spatial or temporal token merging methods Wang et al. (2025c); Zhang & Fu (2025); Li et al. (2025c); Jiang et al. (2025a); Li et al. (2025a); Shao et al. (2025); Song et al. (2024); Chai et al. (2024) progressively discard redundant content, while question-/task-aware strategies Jiang et al. (2025b); Dong et al. (2025); Yao et al. (2025); Song et al. (2025b) tailor retained tokens to the query. These approaches substantially lower FLOPs but still rely on dense attention once tokens are merged. Beyond pure self-attention, Mamba-based or hybrid architectures Jiang et al. (2025a); Ren et al. (2025); Xu et al. (2025b) inject state-space or recurrent modules to approach linear-time inference while preserving long-range dependencies. Also, there exists approach to design data efficient systems for further fine-tuning Li et al. (2025b). Another direction targets the key–value cache during decoding via task-aware sparsification and streaming-friendly memory Qin et al. (2025a); Ning et al. (2025); Kim et al. (2025); Yang et al. (2025e) reduce memory and improve throughput, yet prefill still scales quadratically with sequence length. In contrast to methods that mostly decide where to drop or compress tokens, our approach systematically probe the effectiveness of native sparse attention Yuan et al. (2025a) that restructures attention itself to be learnable and sparse from the ground up. VideoNSA attains near-linear scalability up to 128K tokens and processes over 10,000 frames on a single GPU, outperforming compression-only pipelines on long-video understanding, temporal reasoning, and spatial understanding tasks.
@@ A.2 Sparse Attention Mechanism
Sparse attention is a central strategy for efficient long-context modeling in language and multimodal systems. Surveys (Zhang et al., 2025e) categorize approaches into pattern-based vs. dynamic/learned. Pattern-based sparsity. Methods such as Longformer (Beltagy et al., 2020), StreamingLLM (Xiao et al., 2024), and TriangleMix (He et al., 2025) prescribe fixed local/strided patterns that can be applied training-free; recent multimodal works (Zhang et al., 2025d; Yang et al., 2025b) follow similar principles, while hardware-efficient kernels like Flash Sparse Attention (Yan et al., 2025) further reduce prefill latency. InfLLM-V2 Zhao et al. (2025) uses switchable dense sparse attention to smoothly adapt models from short to long sequences while maintaining consistency and achieving efficient acceleration with high performance. ProxyAttn Wang et al. (2025b) uses representative heads for fine-grained block importance estimation, enabling faster sparse attention with minimal performance loss. Dynamic and trainable sparsity. Content- or gradient-adaptive mechanisms select important connections (e.g., diagonal selection (Tyagi et al., 2025) or lag-relative strategies (Liang et al., 2025)); trainable sparse attention improves long-context reasoning (Gao et al., 2025; Vasylenko et al., 2025; Gao et al., 2024), diffusion-based video generation (Zhang et al., 2025g), and state-space models (Zhan et al., 2025). SLA Zhang et al. (2025f) decomposes attention weights into critical, marginal, and negligible parts, combining sparse and low-rank acceleration to greatly reduce computation while preserving generation quality. Hybrid approaches such as RocketKV (Behnam et al., 2025) combine token/cache compression with learned sparsity, and MMInference (Li et al., 2025d) accelerates modality-aware sparse prefill for VLMs. Despite these advances, most techniques are optimized for text or short multimodal contexts and do not directly address the ultra-long, highly redundant spatio-temporal structure of videos. VideoNSA unifies block-wise compression, salient block selection, and a sliding-window branch under learnable gates that dynamically allocate computation across three native sparse branches (Yuan et al., 2025a). This end-to-end, data-driven design preserves critical global/local dependencies while scaling nearly linearly in both time and memory.
@@ Appendix B Detailed Training Settings
Training hyperparameters for VideoNSA are shown in Table 4. We filter a subset of LLaVA-Video-178K Zhang et al. (2024d) as the training data. For each video, we uniformly sample at 4 frames per second and retain only those with 350–550 frames, resulting in 216K video question–answer pairs from the original 961K pairs in LLaVA-Video-178K Zhang et al. (2024d).
@@ Appendix D More Results on Long-form Video Benchmarks
We take LongVideoBench Wu et al. (2024), LongTimeScope Zohar et al. (2025), MLVU Zhou et al. (2024), and TimeScope Zohar et al. (2025) as representative long-video benchmarks and compare against existing token compression and sparse attention methods. As shown in Table 6, Table 7, Table 8, and Table 9, VideoNSA achieves comparable performance without specialized designs. Moreover, we observe that VideoNSA significantly outperforms the baselines on subtasks related to temporal reasoning and on videos of extended length.
@@ Appendix E More Results on Temporal Reasoning Benchmarks
We take Tomato Shangguan et al. (2024) as the representative temporal reasoning benchmark and compare against existing token compression and sparse attention methods. As shown in Table 10, VideoNSAachieves comparable performance without specialized designs. Moreover, we observe that VideoNSA significantly outperforms the baselines on subtasks including object counting, shape description, and human actions.
@@ Appendix F More Results on Spatial Understanding Benchmarks
We take VSIBench Yang et al. (2025a) as the representative spatial understanding benchmark and compare against existing token compression and sparse attention methods. As shown in Table 11, VideoNSA achieves comparable performance without specialized designs. Moreover, we observe that VideoNSA significantly outperforms the baselines on subtasks including object relative direction, route planning, and object size estimation.
@@
Appendix G Visualization of Attention Pattern in Each Branch
@@
Appendix H More Results on Branch Combination
In this section, we report detailed results of different branch combinations across three domains, including long video understanding (Table 12, Tavke 13, Table 14, and Table 15), temporal reasoning (Table 16), and spatial understanding (Table 17). The corresponding performances are summarized in the table, which highlights how the use of individual branches or their combinations affects downstream tasks.
@@ Appendix H More Results on Branch Combination
Appendix I More results on information scaling study
@@
Table 20: Ablation study results on information scaling of LongVideoBench Wu et al. (2024). Metrics include overall accuracy and task-specific scores across different steps. # TPF stands for token per frame, and # F stands for sampling frame number.
@@
Appendix J More results on attention scaling study
@@
Table 26: LongVideoBench Wu et al. (2024) results across different attention budget strategy. Metrics include overall accuracy and task-specific scores across different steps.
@@
Appendix K Full Gate Values Distribution
@@
Appendix L More Inter-head Gate Similarites Visualization
@@
Appendix M More Analysis about Attention Sinks on Various Sparse Attention Settings
@@
Figure 17(a) indicates that in the compression branch, smaller blocks produce sharper and higher sink peaks at the sequence start, while larger blocks used in training reduce the initial peak but introduce broader low-density diffusion with periodic boundary spikes. The selection sinks in 17(b) remain at consistently low densities under different configurations, suggesting that the top-k filtering mechanism robustly suppresses sink formation across different settings. Figure 17 shows the distribution of attention sinks under different sparse attention settings. When varying the window size, sinks are concentrated near the beginning and decay rapidly with position. Overall, larger windows yield lower sink density but broader coverage, while the training configuration ($w=256$) strikes a middle ground and exhibits sparse periodic clusters in the mid-to-late sequence, reflecting sensitivity to local boundaries learned during training.
Appendix N Dense Attention Sink Visualization

### v2

@@ 9 The Use of Large Language Models
  • Results on additional video understanding benchmarks in Section G.
  • Visualization of attention pattern in each branch in Section H.
  • More results on branch combination in Section I.
  • More results on information scaling study in Section J.
  • Additional context-length scaling results of Qwen2.5-VL in Section K.
  • More results on attention scaling study in Section L.
  • Theoretical analysis of context length and attention budget scaling in Section M.
  • Full gate values distribution in Section N.
  • More inter-head gate similarites visualization in Section O.
  • Benchmark-level gating analysis and PCA visualization in Section P.
  • Additional analysis of training and inference efficiency in Section Q.
  • Additional analysis on CMP latency bottleneck in Section R.
  • More analysis about attention sinks on various sparse attention settings can be found in Section S.
  • Comparison of sparsity patterns between text-only NSA and VideoNSA in Section T.
  • Visualization of attention sinks in dense attention in Section U.
@@ A.1 Efficient Video Understanding
Video understanding systems typically convert videos into long sequences of vision tokens, which can easily exceed GPU memory and slow down inference as the video length grows. To address this, existing work mainly address this by token compression, alternative sequence modeling, and KV-cache compression. One important line of work emphasizes token compression. Spatial or temporal token merging methods Wang et al. (2025c); Zhang & Fu (2025); Li et al. (2025f); Jiang et al. (2025a); Li et al. (2025c); Shao et al. (2025); Song et al. (2024); Chai et al. (2024); Liao et al. (2025); Wu et al. (2025b; 2024b; a; 2024c); Cao et al. (2024); Li et al. (2025a; e); Yang et al. (2023) progressively discard redundant content, while question-/task-aware strategies Jiang et al. (2025b); Dong et al. (2025); Yao et al. (2025); Song et al. (2025b) tailor retained tokens to the query. These approaches substantially lower FLOPs but still rely on dense attention once tokens are merged. Beyond pure self-attention, Mamba-based or hybrid architectures Jiang et al. (2025a); Ren et al. (2025); Xu et al. (2025b); Tao et al. (2025) inject state-space or recurrent modules to approach linear-time inference while preserving long-range dependencies. Also, there exists approach to design data efficient systems for further fine-tuning Li et al. (2025d). Another direction targets the key–value cache during decoding via task-aware sparsification and streaming-friendly memory Qin et al. (2025a); Ning et al. (2025); Kim et al. (2025); Yang et al. (2025e); Liu et al. (2025) reduce memory and improve throughput, yet prefill still scales quadratically with sequence length. In contrast to methods that mostly decide where to drop or compress tokens, our approach systematically probe the effectiveness of native sparse attention Yuan et al. (2025b) that restructures attention itself to be learnable and sparse from the ground up. VideoNSA attains near-linear scalability up to 128K tokens and processes over 10,000 frames on a single GPU, outperforming compression-only pipelines on long-video understanding, temporal reasoning, and spatial understanding tasks.
@@ A.2 Sparse Attention Mechanism
Sparse attention is a central strategy for efficient long-context modeling in language and multimodal systems. Surveys (Zhang et al., 2025e) categorize approaches into pattern-based vs. dynamic/learned. Pattern-based sparsity. Methods such as Longformer (Beltagy et al., 2020), StreamingLLM (Xiao et al., 2024), and TriangleMix (He et al., 2025) prescribe fixed local/strided patterns that can be applied training-free; recent multimodal works (Zhang et al., 2025d; Yang et al., 2025b) follow similar principles, while hardware-efficient kernels like Flash Sparse Attention (Yan et al., 2025) further reduce prefill latency. InfLLM-V2 Zhao et al. (2025) uses switchable dense sparse attention to smoothly adapt models from short to long sequences while maintaining consistency and achieving efficient acceleration with high performance. ProxyAttn Wang et al. (2025b) uses representative heads for fine-grained block importance estimation, enabling faster sparse attention with minimal performance loss. Dynamic and trainable sparsity. Content- or gradient-adaptive mechanisms select important connections (e.g., diagonal selection (Tyagi et al., 2025) or lag-relative strategies (Liang et al., 2025)); trainable sparse attention improves long-context reasoning (Gao et al., 2025; Vasylenko et al., 2025; Gao et al., 2024), diffusion-based video generation (Zhang et al., 2025g), and state-space models (Zhan et al., 2025). SLA Zhang et al. (2025f) decomposes attention weights into critical, marginal, and negligible parts, combining sparse and low-rank acceleration to greatly reduce computation while preserving generation quality. Hybrid approaches such as RocketKV (Behnam et al., 2025) combine token/cache compression with learned sparsity, and MMInference (Li et al., 2025g) accelerates modality-aware sparse prefill for VLMs. Despite these advances, most techniques are optimized for text or short multimodal contexts and do not directly address the ultra-long, highly redundant spatio-temporal structure of videos. VideoNSA unifies block-wise compression, salient block selection, and a sliding-window branch under learnable gates that dynamically allocate computation across three native sparse branches (Yuan et al., 2025b). This end-to-end, data-driven design preserves critical global/local dependencies while scaling nearly linearly in both time and memory.
@@ Appendix B Detailed Training Settings
Training hyperparameters for VideoNSA are shown in Table 4. We filter a subset of LLaVA-Video-178K Zhang et al. (2024e) as the training data. For each video, we uniformly sample at 4 frames per second and retain only those with 350–550 frames, resulting in 216K video question–answer pairs from the original 961K pairs in LLaVA-Video-178K Zhang et al. (2024e).
@@ Appendix D More Results on Long-form Video Benchmarks
We take LongVideoBench Wu et al. (2024a), LongTimeScope Zohar et al. (2025), MLVU Zhou et al. (2024), and TimeScope Zohar et al. (2025) as representative long-video benchmarks and compare against existing token compression and sparse attention methods. As shown in Table 6, Table 7, Table 8, and Table 9VideoNSA achieves comparable performance without specialized designs. Moreover, we observe that VideoNSA significantly outperforms the baselines on subtasks related to temporal reasoning and on videos of extended length.
@@ Appendix E More Results on Temporal Reasoning Benchmarks
We take Tomato Shangguan et al. (2024) as the representative temporal reasoning benchmark and compare against existing token compression and sparse attention methods. As shown in Table 10VideoNSAachieves comparable performance without specialized designs. Moreover, we observe that VideoNSA significantly outperforms the baselines on subtasks including object counting, shape description, and human actions.
@@ Appendix F More Results on Spatial Understanding Benchmarks
We take VSIBench Yang et al. (2025a) as the representative spatial understanding benchmark and compare against existing token compression and sparse attention methods. As shown in Table 11VideoNSA achieves comparable performance without specialized designs. Moreover, we observe that VideoNSA significantly outperforms the baselines on subtasks including object relative direction, route planning, and object size estimation.
@@
Appendix G Results on Additional Video Understanding Benchmarks
We conduct additional experiments on LSDBench Qu et al. (2025) and VideoEvalPro Ma et al. (2025) to compare VideoNSA and other training-free sparse attention baselines, demonstrating the consistent advantage of VideoNSA in multiple video understanding tasks.
Table 12: Results on LSDBench Qu et al. (2025).
Model Accuracy
LongVA Zhang et al. (2024b) 32.5
LongVila Chen et al. (2024b) 49.8
InternVL2.5 Chen et al. (2024c) 50.1
Qwen2.5-VL-7B Qwen et al. (2025) 52.2
Qwen2.5-VL-7B-SFT 52.5
Sparse Attention Methods
+ Tri-Shape Li et al. (2024c) 49.5
+ MInference Jiang et al. (2024) 49.5
+ FlexPrefill Lai et al. (2025) 52.3
+ XAttention Xu et al. (2025a) 51.3
VideoNSA 55.2
Table 13: Results on VideoEvalPro Ma et al. (2025). HP stands for Holistic Perception, HR stands for Holistic Reasoning, LR stands for Local Reasoning, LP stands for Local Perception.
Model HP HR LR LP Overall
LongVA Zhang et al. (2024b) 20.5 6.8 19.0 9.5 16.5
Video-XL Shu et al. (2025) 22.3 15.0 18.2 10.2 18.6
InternVL2.5 Chen et al. (2024c) 28.8 19.7 21.5 16.7 24.6
Qwen2.5-VL-7B Qwen et al. (2025) 33.9 15.6 24.8 17.8 27.7
Qwen2.5-VL-7B-SFT 34.5 15.8 25.3 18.2 28.3
Sparse Attention Methods
+ Tri-Shape Li et al. (2024c) 34.1 16.3 25.1 20.0 28.4
+ MInference Jiang et al. (2024) 32.3 17.1 27.7 16.7 26.0
+ FlexPrefill Lai et al. (2025) 33.0 15.9 26.3 19.8 28.3
+ XAttention Xu et al. (2025a) 34.5 16.6 25.6 20.5 28.9
VideoNSA 35.4 16.9 26.3 19.1 29.4
[…]
black
Appendix H Visualization of Attention Pattern in Each Branch
@@
Appendix I More Results on Branch Combination
In this section, we report detailed results of different branch combinations across three domains, including long video understanding (Table 14, Tavke 15, Table 16, and Table 17), temporal reasoning (Table 18), and spatial understanding (Table 19). The corresponding performances are summarized in the table, which highlights how the use of individual branches or their combinations affects downstream tasks.
@@ Appendix I More Results on Branch Combination
Appendix J More results on information scaling study
@@
Table 22: Ablation study results on information scaling of LongVideoBench Wu et al. (2024a). Metrics include overall accuracy and task-specific scores across different steps. # TPF stands for token per frame, and # F stands for sampling frame number.
@@
Appendix K Additional Context-length Scaling Results of Qwen2.5-VL
((a))
((b))
((c))
((d))
((e))
Figure 15: Performance comparison of Qwen2.5-VL and VideoNSA under different context lengths.
Table 26: Performance of Qwen2.5-VL 7B under different context lengths.
Context LVB MLVU TimeScope LTS VSIBench
32k 58.1 50.5 80.33 34.4 35.1
64k 59.7 51.2 81.13 40.7 29.7
128k 58.7 48.0 81.00 39.7 29.3
Table 27: Performance of VideoNSA under different context lengths.
Context LVB MLVU TimeScope LTS VSIBench
32k 58.7 50.6 78.5 41.1 36.0
64k 59.4 51.8 81.7 44.4 36.6
128k 60.0 48.7 83.7 42.9 35.7
We include Table 26 and Table 27 to further illustrate the long-context behavior of the base model. Since Qwen2.5-VL 7B Qwen et al. (2025) has a maximum context window of 128k, its modeling ability tends to become less stable when approaching this upper bound. As shown in Figure K, Qwen2.5-VL Qwen et al. (2025) often peaks at 64k and slightly declines at 128k across several benchmarks. In contrast, VideoNSA maintains stable or stronger performance at 128k, demonstrating that the observed $64\text{k}>128\text{k}$ phenomenon arises from backbone limitations rather than the proposed sparse architecture.
Appendix L More results on attention scaling study
@@
Table 30: LongVideoBench Wu et al. (2024a) results across different attention budget strategy. Metrics include overall accuracy and task-specific scores across different steps.
@@
Appendix M Theoretical Foundations of Scaling Behavior
In Section 4, we perform two scaling experiments along context length and attention budget. We observe that VideoNSA exhibits strong extrapolation ability on context length: although trained with only 36K tokens, it can generalize to 128K at test time, achieving the best performance at 64K. In contrast, when scaling the attention budget, even a small reduction to 3̃.6% of attention computation already delivers outstanding performance, and further increasing the visible-token count does not yield additional gains. To clarify these phenomena, we provide theoretical interpretations from routing-path stability and the geometric structure of RoPE Su et al. (2024).
Routing-path Stability.
Recent work Huang et al. (2025) indicates that a model’s ability to maintain performance on long sequences depends critically on the stability of its attention routing structure across positions. In the standard attention mechanism, the attention weight from the query vector $Q_{n}$ at position $n$ to the key vector $Z_{j}$ at position $j$ is defined as
$\text{Attn}_{n\to j}=\frac{\exp(Z_{j}^{\top}Q_{n})}{\sum_{k}\exp(Z_{k}^{\top}Q_{n})}.$
Here, $Q_{n}$ and $Z_{j}$ denote the query and key representations at positions $n$ and $j$, respectively. If the model can consistently focus its attention on the task-relevant target token set $\mathcal{T}$ during inference, then (i) $\sum_{j\in\mathcal{T}}\text{Attn}_{n\to j}$ should dominate across different positions; and (ii) the attention assigned to the same key token $j$ should remain nearly unchanged under positional shifts, i.e.,
$\Delta_{i}=\bigl|\text{Attn}_{n\to j}-\text{Attn}_{n+i\to j}\bigr|\approx 0,$
where $\Delta_{i}$ measures the deviation of routing paths across positions in long sequences.
When we scale the context length using dense temporal and spatial sampling, the sparse-attention pattern and mask structure $M$ remain unchanged, which means the model continues to use the routing structure learned during training while simply facing a larger pool of candidate evidence. Since denser sampling mainly introduces redundant or finer-grained details, the model treats these tokens as auxiliary evidence, leaving the core target tokens and their relative attention weights essentially unchanged. Consequently, the overall routing-path structure is preserved, $\Delta_{i}$ remains small, and the model can maintain or even improve its performance at longer context lengths.
In contrast, attention-budget scaling explicitly modifies the set of visible tokens in the sparse-attention mechanism by replacing the original mask $M$ with a new mask $M^{\prime}$. The effective query becomes
$Q_{\text{eff}}=Q\odot M^{\prime},$
where $\odot$ denotes elementwise multiplication, and the corresponding new attention weight is
$\text{Attn}^{\prime}_{n\to j}\propto\exp\bigl(Z_{j}^{\top}(Q\odot M^{\prime})\bigr).$
Even if the modification from $M$ to $M^{\prime}$ appears small in proportion, it substantially changes the set of candidate evidence accessible to each query and alters the relative logits. First, the newly visible tokens reduce the relative weight allocated to the original key tokens, producing a dilution effect. Second, in video tasks, the added visible tokens often lie in similar visual-semantic clusters as the original key tokens and thus have non-negligible similarity scores $Z_{j^{\prime}}^{\top}Q$. These tokens directly compete with and divert attention away from the originally dominant key tokens. Under the combined influence of these effects, the stable routing-path structure learned during training is overwritten: formerly high-weight key tokens may be diluted or overshadowed by the new candidates, causing $\Delta_{i}$ to increase significantly. As a result, the model is more likely to follow incorrect reasoning paths, leading to degraded performance.
Geometric Rotational of RoPE.
RoPE Su et al. (2024) maps the representation at position $i$ into a rotation in a two-dimensional subspace:
$q^{\prime}(i)=R(i\omega)q,\qquad k^{\prime}(j)=R(j\omega)k,$
where $R(i\omega)$ is a rotation matrix and $\omega$ denotes the frequency parameters. This yields an inner product that depends only on the relative distance between the two positions:
$\langle q^{\prime}(i),k^{\prime}(j)\rangle=\langle R((i-j)\omega)q,k\rangle.$
RoPE Su et al. (2024) therefore establishes a structured geometric correspondence between relative distance and rotation phase. Under this geometry, when the context length is moderately increased (e.g., from 36K to 64K), the model only needs to resolve a larger phase difference $d\omega$; within this range, the growth of the phase still lies in the extrapolation regime covered by the empirical distribution seen during training. As a result, the model can naturally generalize.
LM-Infinite Han et al. (2023) further proves that, in order to distinguish the growing clusters of relative distances $\alpha(n)$, the attention logit must increase monotonically with sequence length:
$\sup_{q,k,d\leq n}|w(q,k,d)|\geq\left(\frac{\alpha(n)}{2}\right)^{1/(2r)}\frac{\varepsilon}{4e},$
where $w(q,k,d)$ denotes the logit at relative distance $d$, and $\alpha(n)$ grows with $n$. This “logit growth” is controlled and beneficial at moderate lengths, expanding the dynamic range of attention and enabling the model to maintain token separability over larger distances and consistenting with the strong performance we observe around 64K.
However, when the effective phase difference $d\omega$ becomes excessively large, the rotation angle may approach or exceed the periodic range of multiple frequency dimensions, giving rise to phase aliasing: tokens that should correspond to distinct relative distances collapse into similar or even indistinguishable phase regions. In such cases, although attention logits continue to grow with length, the high-frequency components of RoPE lose their discriminative resolution, reducing geometric separability among tokens, which aligns with existing analyses Press et al. (2021); Chen et al. (2023) showing the degradation of relative positional encoding at extreme distances.
Appendix N Full Gate Values Distribution
@@
Appendix O More Inter-head Gate Similarites Visualization
@@
Figure 18: More Inter-head Gate Similarites Visualization
@@
Figure 19: More Inter-head Gate Similarites Visualization
@@
Figure 20: More Inter-head Gate Similarites Visualization
@@
Figure 21: More Inter-head Gate Similarites Visualization
@@
Figure 22: More Inter-head Gate Similarites Visualization
@@
Figure 23: More Inter-head Gate Similarites Visualization
@@
Figure 24: More Inter-head Gate Similarites Visualization
Appendix P Benchmark-level Gating Analysis and PCA Visualization
((a))
((b))
((c))
Figure 25: PCA visualization of benchmark-level gate patterns.
In this section, we provide additional evidence that VideoNSA’s routing strategy depends on input video content rather than layer depth alone. We collect the layer–head gate vectors for representative videos from three benchmarks with distinct visual properties (LongTimeScope Zohar et al. (2025) for multi-shot transitions, Tomato Shangguan et al. (2024) for high-frequency motion, and VSIBench Yang et al. (2025a) for complex spatial layouts) and project the gate vectors into a 2D space using PCA.
As shown in Figure 25, the gate patterns form three clearly separated clusters, regardless of whether we use the compression branch, the selection branch, or the sliding-window branch, which indicates that VideoNSA learns benchmark-specific routing strategies conditioned on visual content, rather than following a fixed depth pattern.
Table 34: Ablation on static gates averaged over a 1K training subset.
LongVideoBench MLVUTest TimeScope LongTimeScope Tomato VSIBench
Qwen2.5-VL-7B 58.7 51.2 81.0 40.7 22.6 29.7
VideoNSA+ Static Gate 58.4 (-0.5%) 51.2 (0.0%) 81.2 (+0.2%) 41.5 (+2.0%) 23.7 (+4.8%) 31.8 (+7.1%)
[…]
black
To further isolate the role of input-driven routing, we replace each layer’s gate with a static value averaged from a 1K training subset, forcing the model to depend only on layer depth. As shown in Table 34, the performance drops across all six benchmarks, especially on tasks requiring long-range temporal integration, confirming that dynamic gating is essential.
Appendix Q Additional Analysis of Training and Inference Efficiency
Table 35: Theoretical FLOPs comparison among different attention mechanisms. “VideoNSA (ideal)” denotes the theoretical FLOPs of NSA without query-head padding.
Method FLOPs Relative
[…]
((a))
((b))
Figure 26: Inference efficiency comparison across attention mechanisms.
To complement the efficiency discussion in the main paper, we provide additional analysis of both FLOPs and wall-clock latency across different attention mechanisms and context lengths.
Training Efficiency.
Under identical optimization settings, training VideoNSA requires approximately $4600$ H100 GPU hours, while the dense baseline requires $5280$ H100 GPU hours. This corresponds to $0.87\times$ of the dense baseline, indicating that VideoNSA achieves slightly improved training efficiency despite using a more complex attention mechanism.
Inference Efficiency.
Table 35 presents the theoretical FLOPs of different attention mechanisms. In the ideal case, NSA requires only $2.05$ PFLOPs, which is $0.24\times$ that of Flash Attention, demonstrating the theoretical computational efficiency of the sparse routing structure. However, the actual FLOPs and wall-clock latency of VideoNSA are higher than this ideal value due to implementation constraints in the current NSA kernel. The Qwen2.5-VL 7B Qwen et al. (2025) adopts an unusual head configuration of $4$ KV heads and $28$ query heads. To satisfy Triton kernel requirements, the query heads must be padded to $64$, which introduces additional computation and memory access overhead. As a result, the practical efficiency of VideoNSA deviates from its theoretical FLOPs advantage. As shown in Figure 26VideoNSA’s latency grows much more slowly than dense attention, and compared with other sparse baselines, it delivers competitive inference speed while achieving stronger model performance.
Appendix R Additional Analysis on CMP Latency Bottleneck
Table 36: CMP latency under different block sizes and context lengths.
Block Size 4k 8k 16k 32k 64k 128k
32 0.868 1.022 1.311 1.982 3.698 8.343
64 0.882 1.036 1.322 1.997 3.687 8.323
128 0.880 1.027 1.308 1.993 3.705 8.353
Table 37: Latency comparison between different NSA implementations at 8k context length.
Implementation forward backward
nsa-impl Pai et al. (2025a) 5.402 32.826
flash-nsa mdy666 (2025) 2.429 5.537
In this section, we provide additional analysis supporting the observation in findings that the CMP branch becomes the dominant source of latency as the context length increases.
Since the block size determines how many CMP operations are executed, we vary the block size and measure the resulting latency across multiple context lengths. As summarized in Table 36, although increasing the block size reduces the number of CMP executions, the overall latency improvement remains small. Smaller blocks are dominated by memory-access overhead, while larger blocks incur higher computation per block. As a result, block-size scaling affects latency only moderately within a narrow range and does not change the overall scaling trend.
We also observe that the most significant acceleration comes from more efficient NSA implementations instead of architectural hyperparameters. As shown in Table 37, the flash-nsa mdy666 (2025) implementation runs about twice as fast as our current nsa-impl Pai et al. (2025a) in the forward pass and up to six times faster in the backward pass. Other teams are also developing improved kernels such as optimizing NSA for TPUs Ko (2025). These findings show that the dominant factor affecting CMP and overall NSA latency comes from kernel efficiency, including memory access patterns and kernel design.
Appendix S More Analysis about Attention Sinks on Various Sparse Attention Settings
@@
Figure 27(a) indicates that in the compression branch, smaller blocks produce sharper and higher sink peaks at the sequence start, while larger blocks used in training reduce the initial peak but introduce broader low-density diffusion with periodic boundary spikes. The selection sinks in 27(b) remain at consistently low densities under different configurations, suggesting that the top-k filtering mechanism robustly suppresses sink formation across different settings. Figure 27 shows the distribution of attention sinks under different sparse attention settings. When varying the window size, sinks are concentrated near the beginning and decay rapidly with position. Overall, larger windows yield lower sink density but broader coverage, while the training configuration ($w=256$) strikes a middle ground and exhibits sparse periodic clusters in the mid-to-late sequence, reflecting sensitivity to local boundaries learned during training.
Appendix T Discussion on Modality-Specific Sparsity
Figure 28: Layer-wise gate distributions of text-only NSA Pai et al. (2025a).
To further contextualize the modality-specific sparsity patterns exhibited by VideoNSA, we compare its behavior with the text-only NSA Pai et al. (2025a) used in language models. As shown in Figure 28, the text-only NSA Pai et al. (2025a) displays a distinct gating dynamic. The sliding-window branch gradually becomes dominant in deeper layers, while the compression and selection branches diminish rapidly and remain almost inactive throughout most of the network. This pattern reflects the one-dimensional and relatively uniform nature of textual sequences, where long-range interactions are sparse and stable, and the model tends to converge toward a single prevailing routing path. The text-only NSA Pai et al. (2025a) also presents a noticeable anomaly in the final layer, where all three branches suddenly become active again despite having remained largely inactive in previous layers. This behavior suggests a late-stage shift in inductive patterns that is characteristic of language modeling. In contrast, VideoNSA, as shown in Figure 17, maintains active and balanced usage of all three branches across nearly the entire depth of the network, with the compression branch playing a consistently prominent role.
Figure 29: Inter-head gate similarities of text-only NSA Pai et al. (2025a).
The inter-head similarities further highlight the divergence between the two modalities. The text-only NSA Pai et al. (2025a) in Figure 29 exhibits strong correlations across heads in the early layers, which indicates a set of conserved induction-like operations. Later in the model, selection and sliding window gate values become decorrelated across heads. VideoNSA, as shown in Section O, however, displays substantially weaker cross-head correlations overall, and only a few mid-layer clusters emerge in the selection and sliding-window branches. These findings imply that VideoNSA adjusts its sparse routing behavior to accommodate the rich spatiotemporal redundancy and multi-scale structure of video inputs, rather than collapsing into a single dominant pathway as observed in the text-only model.
Appendix U Dense Attention Sink Visualization